Tom,Harry and joan shared sh.29000 among themselves.Tom got sh. 300 more than Harry.Harry got sh.5000 less than Joan.How much did each of them get?
Answer by mananth(16946)


Tom,Harry and joan shared sh.29000 among themselves..How much did each of them get?
let Joan get $x
Harry got sh.5000 less than Joan
x-5000
Tom got sh. 300 more than Harry.
x-5000+300 = x-4700
x+x-5000+x-4700 = 29000
3x-9700 = 29000
3x= 38700
x=12900
Joan gets $12900
You continue